Project

General

Profile

NDOS Commissioning Meeting 2011/09/26

Attendance: Jon, Enrique, Mat, Mark Mathis, Andrew, Jarek, Tricia, Gary, Susan, Minerba, Nick, Mayly, Rafael, Peter, Mark Messier

Agenda:

  1. APD Redistribution Plans
  2. Beam & Detector Update
  3. DAQ Status & Plans
  4. DCS Status & Plans
  5. AOB

APD Redistribution Plans

  • Minerba's studys
    • presented some studies of Cana performance for essentially two APD dist. configurations:
      • "configuration 3": proposed by Ken Heller which sacrifices some vertical APDs in the fiducial volume
      • "configuration 4": proposed by Jon Paley (configuration 4) with input from the physics conveners which sacrifices some fiducial volume for a fully instrumented muon catcher.
      • (4) shows an increase in the # NC events found by 30-40% over (3)
      • (3) shows an increase in the # nu_mu CC events found by ~10% over (4)
      • estimate of # of nu_e CC events with (4) is ~27. Assuming NC/nu_e CC efficiencies are roughly the same, then (3) should give ~20 nu_e CC events.
      • using a data file with a fully instrumented MC, (4) was shown to have a much higher cosmic background rate, but there were several unexplained inconsistencies and it was generally felt that there was something wrong with these results. The cosmic background study will be redone using MC instead of data.
      • "fully contained" events seem to come from muons below 1 GeV, and so Minerba argues that we should have more planes instrumented in DB 3 to improve our energy resolution.
      • another configuration was presented which would reduce the number of APDs that need to be moved from ~93 to 40, but this configuration does not give as complete coverage in the MC, diblock 3 or the fiducial volume
  • Jarek: regardless of final decision, should start moving some APDs soon
  • Jon: proposal is to move APDs from first block (Mat says there are 24) downstream to fiducial volume. Mark: should instrument the vertical planes.
  • Summary:
    • Mark stated a preference for "configuration 4"
    • there were some proponents for configuration 3,
    • Mark and Gary will discuss the plan with the EC and will make a decision soon
    • The first group of ~20 APDs will be moved this week (most likely Wed.-Thurs.)

Beam & Detector

  • This morning the shifter discovered that ~8 APDs appear to be running in a cooled state. This was a false alarm, in fact these channels have no TECCs. The DCS group will consider options for properly displaying the "state" of the APD in the GUI.
  • Over the weekend, the DAQ crashed. This was due to the DDS memory leak which brought the DAQ "master" machine down. Since this machine is the NIS server, logins to all other machines was disabled. The shifter called the DAQ on-call expert, but the expert claimed to have never received (missed) a call, and no voice mail was left on his phone. Perhaps the shifter dialed the wrong number? The situation was eventually resolved Sunday morning.

DAQ Status & Plans

  • Leon: thinks 20 APDs will be ready to install on NDOS. ~Half are recovered, the other half are recovered and paralene-coated
  • Peter: some updates, Jon would like to try to roll out latest version of RC (and Resource Manager); Ganglia free memory plot on "master" to be added to checklist

DCS Status & Plans

  • nothing beyond what was already discussed.

AOB

  • Mat (to Leon):
    • what's the status of the silicone coated APD? Leon: should have already arrived
    • what's the plan of the new mounting hardware? Leon: new parts will be available ~end of October.